OpenStudy (anonymous):

Can anyone please help and explain? Find the distance between the points (-3,2) and (-7,4)

OpenStudy (anonymous):

I know I have to use the Distance Formula, but the answer I am getting does not match the choices I have. The choices are A)\[2\sqrt{10}\] B) \[2\sqrt{5}\] C) 20 D) \[2\sqrt{26}\]